# Key-Point-Driven Mathematical Reasoning Distillation of Large Language Model

### v3

@@ 5 Conclusion
In this paper, we propose Key-Point-Driven Distillation (KPDD) for enhancing mathematical reasoning in Small Language Models (SLMs). Our approach leverages the extraction of key points from questions to improve understanding and reduce errors in reasoning tasks. Experimental results demonstrate that KPDD significantly reduces understanding errors compared to conventional mathematical reasoning distillation method. However, PoTD implicitly embeds the reasoning process within the generated program, making it difficult to analyze misunderstandings. In the future, we will explore error analysis methods to facilitate PoTD error analysis.

### v4

@@ 5 Conclusion
In this paper, we propose Key-Point-Driven Distillation (KPDD) for enhancing mathematical reasoning in Small Language Models (SLMs). Our approach leverages the extraction of key points from questions to improve understanding and reduce errors in reasoning tasks. Experimental results demonstrate that KPDD significantly reduces understanding errors compared to conventional mathematical reasoning distillation method. However, PoTD inherently integrates the reasoning process into the generated program, which complicates error analysis. In the future, we will investigate additional methods for error analysis to enhance the evaluation of PoTD’s performance.
Appendix A Data Generation for KPDD-PoT
Similar to KPDD-CoT, we first sample $k$ examples from a mathematical dataset $\mathcal{D}$, then use these samples to manually create a demonstration set $\mathcal{D}_{p}$. Each reasoning process in the demonstration set $\mathcal{D}_{p}$ includes a core question, problem-solving information, and rationales in PoT format. By utilizing the demonstration set $\mathcal{D}_{p}$ and the instruction, we prompt LLMs to generate reasoning processes in for questions. Figure 5 shows the data synthesis process for KPDD-PoT, and the KPDD-PoT dataset generation process can be formalized as:
$p_{i}=f_{\mathcal{M}}(x_{i},\mathcal{D}_{p}).$ (4)
Appendix B Fine-tuning SLMs for KPDD-PoT
In KPDD-PoT, aside from replacing the KPDD-CoT dataset with the KPDD-PoT dataset, the fine-tuning method for KPDD-PoT-Key remains consistent with that of KPDD-CoT-Key. However, the fine-tuning method of KPDD-PoT-Solve is different with KPDD-CoT-Solve. The main difference between them is the input instruction. Specifically, when fine-tuning KPDD-PoT-solve, the input instruction is "Let’s understand the core question and the problem-solving information, and generate a python program to solve the question." This instruction guides the model to not only understand the core question and the problem-solving information but also to generate Python code that can get the final answer. This approach leverages the model’s ability of code generation, which can be particularly effective for solving mathematical problems programmatically.